Linear functions
y=3/4x + 2, find x if y=2/3

Answer by ijeoma(3)

if y=2/3
plug in the value of y in the y=3/4x and solve for x
therefore; 2/3=3/4x
multiply both sides by 4/3
2/3 x 4/3=3/4(4/3)x
8/9=x
therefore x is 8/9
Proof
substitute the value of x in y=3/4x and u will have the original value of y